# Dark Branches of Immortal Stars at the Galactic Center

### v1

@@ Abstract
In this letter we produce, for the first time, a population-level description of main sequence evolution under the influence of significant dark matter energy injection – and construct a novel dark main sequence. We discover a new branch of higher-mass dark matter powered stars that are observationally distinct from the standard main sequence. The stability of these stars is analogous to the hydrostatic equilibrium of standard stars – as the star expands the dark matter capture rate can decrease (see App. I), causing the stars to contract again. The continuous accretion of dark matter fuel allows these stars to maintain equilibrium eternally, causing the dark main sequence to dominate the stellar population.
@@
Stellar Population Inputs.— To construct our standard and dark HR diagrams, we generate a mock stellar population that evolves both with and without the influence of dark matter. Throughout this letter, we choose relatively standard star formation parameters which are also motivated by the observation of S-cluster stars observed near the Galactic Center [17, 27].
@@
We set the stellar distances to follow a radial distribution that falls as $r^{-2}$ [28, 15, 29], ranging from ${4\times 10^{-5}}$ to ${4\times 10^{-2}}$ pc. Motivated by the migration scenario, where stars cannot form closely to the inner black hole [16], our stellar evolution models in dark matter scenarios do not include any dark matter energy injection until stars enter the main sequence, meaning that the radial binning of stars is irrelevant during their pre-main sequence life. We finally assume that star formation and migration occurs in steady state throughout the 10 Gyr lifetime of our simulations, matching the age of the Milky Way.
@@
Dark Matter Inputs.— We generate two dark matter model benchmarks. Qualitatively, we find similar behavior across both, and show that similar behavior is expected for a wide range of dark matter parameters. In both cases, we assume a dark matter mass of 1 GeV, but vary the following parameters:
Benchmark 1.
We set the dark matter-nucleon scattering cross section to ${\sigma_{\chi N}=3\times 10^{-35}}$ cm2. For this, we follow our constraints on the scattering cross section from Ref. [13] for the star S2, which stands as the best-measured S-cluster object. To set the dark matter density along the different orbits, we choose a generalized NFW profile [30] with index ${\gamma=1.5}$. This corresponds to dark matter densities of $\sim 10^{9}-10^{13}$ GeV/cm3 for the distances considered here.
Benchmark 2.
We choose a smaller dark matter-nucleon scattering cross section of ${\sigma_{\chi N}=1\times 10^{-37}}$ cm2. For the dark matter density, we choose a standard NFW (${\gamma=1.0}$), but include a density spike that arises from the adiabatic accretion of matter by Sagittarius A* [31, 32]. We follow the spike model from [33, 13], $i.e.$ ${\gamma=1.0}$, ${R_{\text{spike}}=10}$ pc and ${\gamma_{\text{spike}}=7/3}$. This results in higher dark matter densities compared to Benchmark 1, and the density increases more steeply towards the Galactic Center, i.e., $\sim 10^{10}-10^{16}$ GeV/cm3.
Our benchmarks are not excluded by direct detection for spin-dependent dark matter models [34, 35, 36, 37, 38], and even given improved direct detection limits in the future, there are a wide range of additional parameter points that provide qualitatively similar results to those we show here. For example, our qualitative results are not affected for lower dark matter masses. In addition, at lower cross sections, large dark matter densities can compensate for any decrease in the signal. Finally, we note that for extremely high densities, the efficient annihilation of dark matter may cause the dark matter density to saturate at a value that depends on the dark matter mass and annihilation cross section. However, in the case of models that use stellar capture to focus dark matter annihilation, our scattering cross-section limits continue to apply even for extremely small annihilation cross-sections that will not saturate up to very large densities.
Stellar Evolution Simulations.— We use the stellar evolution code MESA [39, 40, 41, 42, 43, 44], version r24.03.1 with Software Development Kit (SDK) version 23.7.3, to investigate the effect of dark matter energy injection on a wide range of stellar masses. We base our model on the default work model, but implement a custom module to calculate the capture rate at each time step according to the current stellar radius. To obtain the capture rate, we follow the formalism of Ref. [45], under the assumption of capture-annihilation equilibrium, and add the annihilation energy into the stellar core through the inject_uniform_extra_heat parameter. The injection of the additional energy can be restricted to a certain mass region of the star through the parameters min_q_for_uniform_extra_heat and max_q_for_uniform_extra_heat. We consider the stellar core to correspond to the inner 10% of the star, but have tested that injecting the dark matter annihilation energy in the inner 1% only results in slight numerical differences, while uniform injection throughout the star only changes our results by a factor of a few, which does not have qualitative effects. We assume that all stars initially form and evolve in a low dark matter density environment until their early main sequence phase, $i.e.$ their pre-main sequence evolution is not affected by dark matter. We assume that these stars migrate as main sequence stars towards the Galactic Center where their subsequent evolution is influenced by dark matter annihilation. As discussed above, this migration scenario is a commonly considered case, due to the fact that stellar formation close to Sagittarius A* is expected to be inhibited [16]. In MESA, we start injecting the additional energy when the hydrogen fraction in the star center drops below 0.67, which corresponds to an early main sequence core composition, through the parameter xa_central_lower_limit. Once the star experiences dark matter energy injection, we evolve the star until it reaches the red giant branch phase, using the stopping condition stop_at_phase_ZACHeB, or until an age of 10 Gyr is reached, corresponding to the age of the Galaxy.
Results and Discussion.— Figure 2 shows the effects of dark matter energy injection on stars of different masses and distances from the Galactic Center for Benchmark 1 and 2. These models only differ in the dark matter-nucleon scattering cross section and the underlying dark matter profile, that determines the amount of dark matter a star can capture along its orbit. We find that stellar evolution is affected in various ways:
@@
    Stuck on Hayashi track. The orange region corresponds to dark matter powered stars that are found in the region of the Hayashi track in the HR diagram, which, in the standard picture, describes pre-main sequence stars that are still contracting and have not yet started nuclear fusion. At the onset of dark matter burning, main sequence stars evolve “backwards”, similar to the stars halting on the Henyey track described above. In this scenario however, dark matter burning replaces nuclear fusion entirely, and stars move all the way to the Hayashi track. The stars do not undergo any further evolutionary stages as long as the dark matter capture rate remains sufficiently high to prevent nuclear fusion from resuming. A star like this might look like a young, still-forming star but has features of a star that has undergone nuclear fusion in the past – and is effectively immortal.
@@
    Beyond Hayashi track. The dark orange region corresponds to stars moving back onto the Hayashi track, similar to case (3). However, in this case, the stars move further up the Hayashi track to even earlier evolutionary states. In our MESA simulations, these stars end up higher on the Hayashi track than their starting position. We emphasize that this scenario is highly reliant on the modelling methods (as MESA does not simulate of the very early stages of star formation). It is possible that these stars could expand even further or be entirely disrupted by the dark matter energy injection. A scenario like this may be consistent with G-objects, a recently discovered strange class of objects found at similar orbits around Sgr A* as the S-cluster stars that resemble clouds of gas surrounding a denser core [46, 47].
@@
Figure 3: Dark HR diagram for a population of Galactic Center stars, for our dark matter benchmark 2, with 1-GeV dark matter mass, scattering cross section of $10^{-37}$ cm2, and NFW dark matter profile with a spike (see text for details).
Figure 1 shows HR diagrams for a population of Galactic Center stars with and without dark matter for Benchmark 1, and Figure 3 shows the dark HR diagram for the our additional Benchmark 2. We include evolutionary stages from the early main sequence up until the stars leave the main sequence and start hydrogen shell burning. We do not include stars that have become red giants and start core Helium burning, $i.e.$ most of the stars on the HR diagram are main sequence stars. Most stars lie on the main sequence, $i.e.$ the diagonal line, where more massive stars are at the top left, as they are hotter and more luminous, and lighter stars are at the bottom right. The small number of stars scattered to the right of the main sequence are stars that are older stars that have just moved away from the main sequence towards the red giant branch; which manifests in a lower temperature compared to the main sequence.
@@
Finally, we note that the fraction of high-mass stars increases in dark matter scenarios due to the effect of stellar immortality. If only standard stellar models were considered, the large fraction of high-mass stars would be misconstrued as evidence of a top-heavy initial mass function (IMF). Intriguingly, observations have found evidence of a top-heavy IMFs at the Galactic center [48, 20, 21]. We find that dark matter would harden the IMF, compared to a Salpeter IMF, by approximately ${\sim 0.5-2.5}$, as shown in Figs. 1 and 3.
Summary and Outlook.— We have produced the first population-level prediction for stellar evolution under the presence of dark matter capture and annihilation near the Galactic Center. We have established a new “dark main sequence” that is observable in the Hertzsprung-Russell diagram of stars observed in dark matter rich environments. This sequence has three observable characteristics that clearly separate it from standard main sequence stellar evolution: (1) a top-heavy initial mass function along the main sequence branch, produced by the efficient removal of low-mass stars off of the main sequence and into locations on the Hayashi or Beyond-Hayashi tracks, (2) a bright and densely populated Hayashi track, produced by low-mass main sequence stars which achieve eternally stable equilibria in configurations with large stellar radii and low stellar temperatures, (3) a bright and densely populated Henyey track at higher-masses, populated by effectively immortal (typical lifespans $>$10 Gyr) high-mass stars with similar luminosities but slightly lower temperatures than their main sequence configurations.
Excitingly, there are several anomalies in Galactic Center stellar observations which can be simultaneously explained by stellar evolution along our dark main sequence. The top-heavy initial mass function near the Galactic Center is automatically produced by our simulations with dark matter, while both the paradox of youth and conundrum of old age can be explained if relatively high-mass stars are powered indefinitely by dark matter annihilation. Finally, observations of G-objects (extremely large gas clouds which may hold a central star), may be consistent with Beyond-Hayashi evolution, although the stability of such solutions falls beyond the parameters of our current MESA modeling. We will investigate this in upcoming work.
@@
Acknowledgments.— We thank Andrea Ghez for helpful discussions. This work was supported by a collaborative visit funded and hosted by Stanford KIPAC and the Cosmology and Astroparticle Student and Postdoc Exchange Network (CASPEN). I.J. and T.L. acknowledge support by the Swedish Research Council under contract 2022-04283. T.L. also acknowledges support from the Swedish National Space Agency under contract 117/19. R.K.L. is supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2023/3-21.
Dark Branches of Immortal Stars at the Galactic Center
Supplemental Material
Isabelle John, Rebecca K. Leane, and Tim Linden
@@
I Dark Matter Capture Rates for Actively Evolving Stars
Refer to captionRefer to caption
@@ I Dark Matter Capture Rates for Actively Evolving Stars
Figure S1: Left panel: Dark matter capture rate as a function of stellar radius, for a scattering cross section of $10^{-37}$ cm2. Right panel: Capture rate as a function of scattering cross section, for stellar radii of 1 and 10 solar radii, and a fixed stellar mass of 1 solar mass. The dashed line shows the transition cross sections for the two stars. To calculate these capture rates, we assume an example dark matter mass of 0.1 GeV, orbital velocity of $10^{8}$ cm/s and dark matter density $2\times 10^{12}$ GeV/cm3.
@@
Figure S1 (left panel) shows the dark matter capture rates expected as the star undergoes evolution, $i.e.$ taking into account the changing mass and radius of the star. The first point to note, as highlighted by the left figure, is that the capture rate changes as the stellar radius changes, and so it is important to take the updated stellar radius into account for the capture rate as the star evolves. This implementation is an improvement over previous studies [13]. The second point is that for a fixed stellar mass and fixed dark matter scattering cross section, the dark matter capture rate can decrease as the stellar radius increases. The reason for this can be understood from the right figure.
@@
Figure S1 (right panel) shows the capture rate as a function of scattering cross section, for two example stars of radius 1 and 10 times the solar radius, for a fixed stellar mass of 1 solar mass. We also show as the dashed line the transition cross section for these two example stars. Here, the transition cross section is $\sigma_{\rm tr}=\pi R^{2}/N$, where $N$ is the number of stellar nucleons available for scattering. Physically, this corresponds to a mean free path of the dark matter of about the size of the star, and therefore corresponds to the cross section above which multiple scatters can be needed for capture to occur. Below this cross section, dark matter can only be captured with roughly one scatter, as the scattering is so weak that on average one or zero scatters occur. As the mass of these two example stars stays the same, the number of targets also remains the same, except they are distributed over a larger volume, the transition cross section between the two simply scales as their radius ratio squared.
@@
As Fig. S1 (right) shows, larger radii stars do capture more dark matter in the sufficiently-high cross section regime where roughly all incoming dark matter is captured, as their surface areas available for capture are larger. However, less dense objects have lower escape velocities, so require more scatters (i.e. a larger cross section) to bring the dark matter to a low enough speed to be captured, which is why the larger star needs a larger cross section increase above its transition cross section at $\sim 10^{-33}$ cm2 to reach its geometric capture rate. In comparison, the smaller radius star increases less above its $\sim 10^{-35}$ cm2 transition before reaching geometric capture, because it is more efficient at capture (and so needs fewer scatters) due to its higher escape velocity. Once comparing the capture rates for these two objects at their transition cross section, the puffier star is already down a larger penalty compared to its geometric maximum capture than the more dense star, which is then cancelled out by its size, such that at their respective transition cross sections (which are two orders of magnitude apart) they actually have the same capture rate. This means that, because the capture rate decreases linearly with cross section below the transition cross section, the puffier, larger radius star has a lower capture rate at smaller cross sections, as its transition cross section is higher to start with. This can lead to interesting effects during the stellar evolution process as the star becomes larger, as the dark matter capture and annihilation driving the growth in radii consequently decreases.

### v2

@@ Abstract
In this article we produce, for the first time, a population-level description of main sequence evolution under the influence of significant dark matter energy injection. We construct a novel dark main sequence, informed, for the first time, by cutting-edge observations of the S-cluster stellar population near the Galactic Center [14, 15]. Specifically, we discover a new branch of higher-mass dark matter powered stars that are observationally distinct from the standard main sequence. The stability of these stars is analogous to the hydrostatic equilibrium of standard stars – as the star expands the dark matter capture rate can decrease (see App. A), causing the stars to contract again. The continuous accretion of dark matter fuel allows these stars to maintain equilibrium eternally, causing the dark main sequence to dominate the stellar population.
@@
Interestingly, our initial data regarding stars near the Galactic Center already unveils numerous anomalies that might be solved by the dark main sequence. Star formation models suggest that stars cannot form within $\sim$ 0.1 pc of the central black hole, where the S-cluster stars are found [17]. Rather, the stars must have formed elsewhere and migrated towards the Galactic Center [14]. Conversely, observations suggest that stars in this region are young ($\lesssim 15$ Myr), which indicates that the stars might have formed more locally [18]. The tension between these two observations is known as the “paradox of youth” problem: Galactic Center stars are as bright as young stars but show spectroscopic features of more evolved stars [14, 19]. Additionally, there is a “conundrum of old age”, which describes the paucity of old and evolved stars, despite the fact that models predict the detection of many evolved stars [20, 21]. Finally, observations indicate that the initial mass function of stars near the Galactic Center is surprisingly top-heavy (biased towards high-mass stars), with observations indicating that the stellar birth distribution lies between $M^{-1.7}$ [22] and $M^{-0.45}$ [23], compared to a standard Salpeter index of $M^{-2.35}$ [24].
@@
I Stellar Population Inputs
To construct our standard and dark HR diagrams, we generate a mock stellar population that evolves both with and without the influence of dark matter. We choose relatively standard star formation parameters which are also motivated by the observation of S-cluster stars observed near the Galactic Center [18, 29].
@@ I Stellar Population Inputs
We set the stellar distances to follow a radial distribution that falls as $r^{-2}$ [30, 17, 31], with 43 bins ranging from ${4\times 10^{-5}}$ to ${4\times 10^{-2}}$ pc. Motivated by the migration scenario, where stars cannot form closely to the inner black hole [14], our stellar evolution models in dark matter scenarios do not include any dark matter energy injection until stars enter the main sequence, meaning that the radial binning of stars is irrelevant during their pre-main sequence life. We finally assume that star formation and migration occurs in steady state throughout the 10 Gyr lifetime of our simulations, matching the age of the Milky Way.
@@
II Dark Matter Inputs
We generate two dark matter model benchmarks. Qualitatively, we find similar behavior across both, and show that similar behavior is expected for a wide range of dark matter parameters. In both cases, we assume a dark matter mass of 1 GeV, but vary the following parameters:
Benchmark 1.
We set the dark matter-nucleon scattering cross section to ${\sigma_{\chi N}=3\times 10^{-35}}$ cm2. For this, we follow our constraints on the scattering cross section from Ref. [13] for the star S2, which stands as the best-measured S-cluster object. To set the dark matter density along the different orbits, we choose a generalized NFW profile [32] with index ${\gamma=1.5}$. This corresponds to dark matter densities of $\sim 10^{8}-10^{12}$ GeV/cm3 for the distances considered here.
Benchmark 2.
We choose a smaller dark matter-nucleon scattering cross section of ${\sigma_{\chi N}=1\times 10^{-37}}$ cm2. For the dark matter density, we choose a standard NFW (${\gamma=1.0}$), but include a density spike that arises from the adiabatic accretion of matter by Sagittarius A* [33, 34]. We follow the spike model from [35, 13], $i.e.$ ${\gamma=1.0}$, ${R_{\text{spike}}=10}$ pc and ${\gamma_{\text{spike}}=7/3}$. This results in higher dark matter densities compared to Benchmark 1, and the density increases more steeply towards the Galactic Center, i.e., $\sim 10^{8}-10^{15}$ GeV/cm3.
Our benchmarks are not excluded by direct detection for spin-dependent dark matter models [36, 37, 38, 39, 40], and even given improved direct detection limits in the future, there are a wide range of additional parameter points that provide qualitatively similar results to those we show here. For example, our qualitative results are not affected for lower dark matter masses. In addition, at lower cross sections, large dark matter densities can compensate for any decrease in the signal. Finally, we note that for extremely high densities, the efficient annihilation of dark matter may cause the dark matter density to saturate at a value that depends on the dark matter mass and annihilation cross section. Even incorporating any blunting we find qualitatively new results for both $s$-wave or $p$-wave annihilation model benchmarks, see Appendix B for more details.
III Stellar Evolution Simulations
We use the stellar evolution code MESA [41, 42, 43, 44, 45, 46], version r24.03.1 with Software Development Kit (SDK) version 23.7.3, to investigate the effect of dark matter energy injection on a wide range of stellar masses. We base our model on the default work model, but implement a custom module to calculate the capture rate at each time step according to the current stellar radius. To obtain the capture rate, we follow the formalism of Ref. [47], under the assumption of capture-annihilation equilibrium, and add the annihilation energy into the stellar core through the inject_uniform_extra_heat parameter. The injection of the additional energy can be restricted to a certain mass region of the star through the parameters min_q_for_uniform_extra_heat and max_q_for_uniform_extra_heat. We consider the stellar core to correspond to the inner 10% of the star, but have tested that injecting the dark matter annihilation energy in the inner 1% only results in slight numerical differences, while uniform injection throughout the star only changes our results by a factor of a few, which does not have qualitative effects. We assume that all stars initially form and evolve in a low dark matter density environment until their early main sequence phase, $i.e.$ their pre-main sequence evolution is not affected by dark matter. We assume that these stars migrate as main sequence stars towards the Galactic Center where their subsequent evolution is influenced by dark matter annihilation. As discussed above, this migration scenario is a commonly considered case, due to the fact that stellar formation close to Sagittarius A* is expected to be inhibited [14]. In MESA, we start injecting the additional energy when the hydrogen fraction in the star center drops below 0.67, which corresponds to an early main sequence core composition, through the parameter xa_central_lower_limit. Once the star experiences dark matter energy injection, we evolve the star until it reaches the red giant branch phase, using the stopping condition stop_at_phase_ZACHeB, or until an age of 10 Gyr is reached, corresponding to the age of the Galaxy. We do not include any energy transport effects directly due to internal dark matter-nucleon scattering. In our scenario these effects are subdominant, as the energy rate injected by dark matter burning far outweighs any effects of the energy transport (see e.g. [48]).
@@ III Stellar Evolution Simulations
IV Results and Discussions
Figure 2 shows the effects of dark matter energy injection on stars of different masses and distances from the Galactic Center for Benchmark 1 and 2. These models only differ in the dark matter-nucleon scattering cross section and the underlying dark matter profile, that determines the amount of dark matter a star can capture along its orbit. We find that stellar evolution is affected in various ways:
@@ IV Results and Discussions
    Stuck on Hayashi track. The orange region corresponds to dark matter powered stars that are found in the region of the Hayashi track in the HR diagram, which, in the standard picture, describes pre-main sequence stars that are still contracting and have not yet started nuclear fusion. At the onset of dark matter burning, main sequence stars evolve “backwards”, similar to the stars halting on the Henyey track described above. In this scenario however, dark matter burning replaces nuclear fusion entirely, and stars move all the way to the Hayashi track, consistent with the results in Refs. [10, 11]. The stars do not undergo any further evolutionary stages as long as the dark matter capture rate remains sufficiently high to prevent nuclear fusion from resuming. A star like this might look like a young, still-forming star but has features of a star that has undergone nuclear fusion in the past – and is effectively immortal.
@@
    Beyond Hayashi track. The dark orange region corresponds to stars moving back onto the Hayashi track, similar to case (3). However, in this case, the stars move further up the Hayashi track to even earlier evolutionary states. In our MESA simulations, these stars end up higher on the Hayashi track than their starting position. We emphasize that this scenario is highly reliant on the modelling methods (as MESA does not simulate the very early stages of star formation). It is possible that these stars could expand even further or be entirely disrupted by the dark matter energy injection. A scenario like this may be consistent with G-objects, a recently discovered strange class of objects found at similar orbits around Sgr A* as the S-cluster stars that resemble clouds of gas surrounding a denser core [49, 50].
@@
Figure 3: Dark HR diagram for a population of Galactic Center stars, for our dark matter benchmark 2, with 1 GeV dark matter mass, scattering cross section of $10^{-37}$ cm2, and NFW dark matter profile with a spike (see text for details).
Figure 1 shows HR diagrams for a population of Galactic Center stars with and without dark matter for Benchmark 1, and Figure 3 shows the dark HR diagram for the additional Benchmark 2. We include evolutionary stages from the early main sequence up until the stars leave the main sequence and start hydrogen shell burning. We do not include stars that have become red giants and start core Helium burning, $i.e.$ most of the stars on the HR diagram are main sequence stars. For more details, see Appendix C. Most stars lie on the main sequence, $i.e.$ the diagonal line, where more massive stars are at the top left, as they are hotter and more luminous, and lighter stars are at the bottom right. The small number of stars scattered to the right of the main sequence are stars that are older stars that have just moved away from the main sequence towards the red giant branch; which manifests in a lower temperature compared to the main sequence.
@@
Finally, we note that the fraction of high-mass stars increases in dark matter scenarios due to the effect of stellar immortality. If only standard stellar models were considered, the large fraction of high-mass stars would be misconstrued as evidence of a top-heavy initial mass function (IMF). Intriguingly, observations have found evidence of a top-heavy IMF at the Galactic Center [51, 22, 23]. We find that dark matter would harden the IMF, compared to a Salpeter IMF, by approximately ${\sim 0.5-2.5}$, as shown in Figs. 1 and 3.
V Summary and Outlook
We have produced the first population-level prediction for stellar evolution under the presence of dark matter capture and annihilation near the Galactic Center. We have established a new “dark main sequence” that is observable in the Hertzsprung-Russell diagram of stars observed in dark matter rich environments. This sequence has three observable characteristics that clearly separate it from standard main sequence stellar evolution: (1) a top-heavy initial mass function along the main sequence branch, produced by the efficient removal of low-mass stars off of the main sequence and into locations on the Hayashi or Beyond-Hayashi tracks, (2) a bright and densely populated Hayashi track, produced by low-mass main sequence stars which achieve eternally stable equilibria in configurations with large stellar radii and low stellar temperatures, (3) a bright and densely populated Henyey track at higher-masses, populated by effectively immortal (typical lifespans $>$10 Gyr) high-mass stars with similar luminosities but slightly lower temperatures than their main sequence configurations.
Excitingly, there are several anomalies in Galactic Center stellar observations which can be simultaneously explained by stellar evolution along our dark main sequence. The top-heavy initial mass function near the Galactic Center is automatically produced by our simulations with dark matter, while both the paradox of youth (see also [11]) and conundrum of old age can be explained if relatively high-mass stars are powered indefinitely by dark matter annihilation. Finally, observations of G-objects (extremely large gas clouds which may hold a central star), may be consistent with Beyond-Hayashi evolution, although the stability of such solutions falls beyond the parameters of our current MESA modeling. We will investigate this in upcoming work.
@@ Acknowledgements
We thank Andrea Ghez for helpful discussions. This work was supported by a collaborative visit funded and hosted by Stanford KIPAC and the Cosmology and Astroparticle Student and Postdoc Exchange Network (CASPEN). I.J. and T.L. acknowledge support by the Swedish Research Council under contract 2022-04283. T.L. also acknowledges support from the Swedish National Space Agency under contract 117/19. R.K.L. is supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2023/3-21.
@@
Appendix A Dark Matter Capture Rates for Actively Evolving Stars
@@
Figure 4: Left panel: Dark matter capture rate as a function of stellar radius, for a scattering cross section of $3\times 10^{-35}$ cm2 (Benchmark 1). Right panel: Capture rate as a function of scattering cross section, for stellar radii of 1 and 10 solar radii, and a fixed stellar mass of 1 solar mass. The dashed line shows the transition cross sections for the two stars. To calculate these capture rates, we assume an example dark matter mass of 1 GeV, orbital velocity of $10^{9}$ cm/s and dark matter density $2\times 10^{12}$ GeV/cm3.
@@
Figure 4 (left panel) shows the dark matter capture rates expected as the star undergoes evolution, $i.e.$ taking into account the changing mass and radius of the star. The first point to note, as highlighted by the left figure, is that the capture rate changes as the stellar radius changes, and so it is important to take the updated stellar radius into account for the capture rate as the star evolves. This implementation is an improvement over previous studies [13]. The second point is that for a fixed stellar mass and fixed dark matter scattering cross section, the dark matter capture rate can decrease as the stellar radius increases. The reason for this can be understood from the right figure.
@@
Figure 4 (right panel) shows the capture rate as a function of scattering cross section, for two example stars of radius 1 and 10 times the solar radius, for a fixed stellar mass of 1 solar mass. We also show as the dashed line the transition cross section for these two example stars. Here, the transition cross section is $\sigma_{\rm tr}=\pi R^{2}/N$, where $N$ is the number of stellar nucleons available for scattering. Physically, this corresponds to a mean free path of the dark matter of about the size of the star, and therefore corresponds to the cross section above which multiple scatters can be needed for capture to occur. Below this cross section, dark matter can only be captured with roughly one scatter, as the scattering is so weak that on average one or zero scatters occur. As the mass of these two example stars stays the same, the number of targets also remains the same, except they are distributed over a larger volume, the transition cross section between the two simply scales as their radius ratio squared.
@@
As Fig. 4 (right) shows, larger radii stars do capture more dark matter in the sufficiently-high cross section regime where roughly all incoming dark matter is captured, as their surface areas available for capture are larger. However, less dense objects have lower escape velocities, so require more scatters (i.e. a larger cross section) to bring the dark matter to a low enough speed to be captured, which is why the larger star needs a larger cross section increase above its transition cross section at $\sim 10^{-33}$ cm2 to reach its geometric capture rate. In comparison, the smaller radius star increases less above its $\sim 10^{-35}$ cm2 transition before reaching geometric capture, because it is more efficient at capture (and so needs fewer scatters) due to its higher escape velocity. Once comparing the capture rates for these two objects at their transition cross section, the puffier star is already down a larger penalty compared to its geometric maximum capture than the more dense star, which is then cancelled out by its size, such that at their respective transition cross sections (which are two orders of magnitude apart) they actually have the same capture rate. This means that, because the capture rate decreases linearly with cross section below the transition cross section, the puffier, larger radius star has a lower capture rate at smaller cross sections, as its transition cross section is higher to start with. This can lead to interesting effects during the stellar evolution process as the star becomes larger, as the dark matter capture and annihilation driving the growth in radii consequently decreases.
For our benchmark, we assume all dark matter remains captured, $i.e.$ that it does not evaporate. This is a good approximation for the stars we consider, in the dense Galactic Center environment.
Appendix B Dark Matter Annihilation Rates
B.1 Impact of $s$-wave vs $p$-wave Annihilation on the Dark Matter Density
Depending on the size of the dark matter annihilation rate, the dark matter density can saturate, i.e. not increase further, towards the Galactic Center. In the main text we do not show the impacts of this blunting, for completeness in this section we will now investigate results for two common scenarios, $s$-wave or $p$-wave annihilation, both of which can arise in a broad range of particle models. We will show that the variance in the expected dark-matter density provides an interesting handle on the size of the dark matter annihilation rate, through different stellar evolution behavior on our dark HR diagrams, with striking qualitatively different behavior compared to the standard HR diagram in all cases.
This saturation density due to dark matter annihilation in the Galaxy is given by
$\rho_{\text{sat}}=\frac{m_{\text{DM}}}{\langle\sigma v\rangle T_{\text{BH}}},$ (2)
where $m_{\text{DM}}$ is the mass of the dark matter particle, $\langle\sigma v\rangle$ is the dark matter annihilation cross section, and ${T_{\text{BH}}\sim 10^{10}}$ yr is the age of Sgr A*, which corresponds to the time dark matter has been accumulated at the Galactic Center.
B.1.1 S-wave Annihilation
Figure 5 illustrates an example of the stellar evolution for classes of particle models with $s$-wave annihilation rates ($\langle\sigma v\rangle=3\times 10^{-26}$ cm3/s), at an example mass of 1 GeV, corresponding to a dark matter saturation limit of $\sim 10^{8}$ GeV/cm3. For comparison, the left panel shows the standard HR diagram, while the right panel shows the dark HR diagram. We observe that the dark HR diagram shows striking differences compared to the standard HR diagram. Intermediate mass stars ($\sim 8-13$ $M_{\odot}$), are observed to populate a distinct dark main sequence (formed from stars residing along the Henyey tracks), albeit thinner than in the non-saturated case (see Fig. 1). Even lighter stars are disrupted already at much lower dark matter densities of $\sim 10^{6}$ GeV/cm3, which is already achieved for the standard NFW profile ($\gamma=1$[13]. This means that even at much lower densities than considered in our Benchmark 1 and 2 scenarios (Figs. 1 and 3), the dark HR diagram displays clear differences to the standard scenario, lacking light stars on the main sequence and strongly populating the Hayashi tracks.
Figure 5: Similar to Fig. 1, where the left panel shows the standard main sequence for stars without dark matter burning, while the right panel presents the a case similar to our Benchmark 1 scenario, but with a dark matter density that saturates at $\sim 5\times 10^{8}$ GeV/cm3 towards the Galactic Center, which is applicable to model classes with $s$-wave annihilation. Compared to the un-saturated case in Fig. 1, the dark main sequence is reduced but still clearly visible, and regions of the lighter stars show similar differences as compared to the standard HR diagram.
B.1.2 P-wave Annihilation
In the case of particle model classes with $p$-wave annihilation ($\langle\sigma v\rangle\approx 10^{-31}$ cm3/s), the saturation density in the Galaxy is much higher than $s$-wave cases, due to the lower annihilation rates in the Galaxy (see Eq. 2). Assuming for example an NFW dark matter profile [32] with a profile index of $\gamma=1.5$, the maximum dark matter density, achieved at the innermost location of S-cluster stars, corresponds to $\sim 10^{12}$ GeV/cm3. This would produce the same dark HR diagrams as presented in the main text, as the blunting is then not important at the Galactic radii of the S-cluster stars. While we observe qualitatively different behaviour for both the $s$-wave and $p$-wave results compared to the standard HR diagram, comparing Fig. 1 with Fig. 5, it is also clear that our setup can provide powerful insights into the dark matter density distribution and dark matter self-annihilation rate, through a statistically different number of stars being affected the same way.
B.2 Dark Matter Equilibrium
Throughout our analysis, we assume that the dark matter capture rate and annihilation rate inside the star are in equilibrium. As we assume that stars form in a region of low dark matter density and migrate towards the Galactic Center during their main sequence phase, some time passes until the dark matter capture rate and annihilation rate inside the star reach equilibrium. This equilibrium time scale is estimated by [52]
$\tau_{\text{eq}}=\sqrt{\frac{V_{\star}}{C\langle\sigma v\rangle}},$ (3)
where $V_{\star}=4/3\times\pi R_{\star}^{3}$ is the volume of the star for a stellar radius $R_{\star}$ and $C$ the dark matter capture rate [47, 13]. Thus, the equilibrium time scale depends on the annihilation cross section (as well as on the dark matter density and other parameters through the capture rate).
In the most extreme case we can consider the above derived required annihilation cross section of $\sim 10^{-31}$ cm3/s, for which the equilibrium time scale is $\sim 10^{6}$ yr. The age of the S-cluster stars is not exactly known, but estimated to be in the order of a few to a few hundreds of Myr [15] – however, if dark matter extends the lifetime of the stars, they could be much older. Therefore, we expect our assumption of equilibrium to be valid.
Appendix C Computation of Hertzprung-Russell Diagrams
We now provide our methodology to obtain the HR diagrams shown in Figs. 13 and 5 from our simulated stellar population, as shown in Fig. 2.
In our HR diagrams, we only include main sequence stars, i.e. we do not show evolutionary stages before and after the main sequence (as we defined in MESA). This is because the S-cluster stars are classified as main sequence stars and we want to re-create HR diagrams that would closely resemble the S cluster.
We then populate the HR diagram by selecting stars according to (1) an initial mass distribution following the Salpeter distribution in Eq. 1, (2) a radial distribution of distances from the Galactic Center (corresponding to different dark matter densities, which affects the stars differently) and (3) an age from 0 to 10 billion years. If for a star of a selected age the current evolutionary stage does not correspond to the main sequence, we do not include this star on the HR diagram, i.e. we do not include the star if it has evolved beyond the main sequence or if it is too young to have reached the main sequence.
Note that when obtaining the HR diagram for the standard main sequence (Fig. 1 bottom panel), the radial distance does not matter, since in the standard scenario we assume that stellar evolution is not changed by dark matter.